# Machine Learning with Physics Knowledge for Prediction: A Survey

### v1

@@ Abstract
This survey examines the broad suite of methods and models for combining machine learning with physics knowledge for prediction and forecast, with a focus on partial differential equations. These methods have attracted significant interest due to their potential impact on advancing scientific research and industrial practices by improving predictive models with small- or large-scale datasets and expressive predictive models with useful inductive biases. The survey has two parts. The first considers incorporating physics knowledge on an architectural level through objective functions, structured predictive models, and data augmentation. The second considers data as physics knowledge, which motivates looking at multi-task, meta, and contextual learning as an alternative approach to incorporating physics knowledge in a data-driven fashion. Finally, we also provide an industrial perspective on the application of these methods and a survey of the open-source ecosystem for physics-informed machine learning.
@@ 1 Introduction
Prediction lies at the heart of science and engineering, and many advances involve the discovery of simple patterns — often expressed as symbolic expressions — that approximate the evolution of our physical world. Still, these mathematical models are only achieved through some degree of simplification and abstraction and thus rarely capture the complexity of the physical system in its full fidelity. As a result, there is significant interest in learning models from only measurements of the real world, spurring the development of machine learning (ML) (Bishop and Nasrabadi, 2006) for the physical sciences. This field relies on using large datasets of measurements to constrain highly flexible parametric models rather than evoking domain knowledge.
@@
However, for applications that require a high degree of reliability, robustness, and trust in their predictions, purely data-driven models could be deemed insufficient, especially when there is limited data. A natural compromise is to build predictive models that leverage the prior knowledge obtained through centuries of scientific study with the flexibility and scale of modern machine learning (Karniadakis et al., 2021; Karpatne et al., 2022). Achieving this requires models that can integrate prior knowledge in the form of inductive biases (Baxter, 2000), e.g., additional task-informed structure incorporated into the model, objective, or learning algorithm. This added structure acts like a filter, essentially guiding the model to focus on solutions that align with scientific understanding. As a result, the model’s search space becomes smaller, with less irrelevant territory to explore. This targeted exploration allows the model to learn effectively even with limited data. The key challenge lies in designing these inductive biases carefully. If the model is too restrictive, the resulting solution might be suboptimal due to underfitting. This review surveys such methods to inform machine learning models with prior physics knowledge for prediction.
Section Solve Infer Generalize Inductive bias
3.1: Differential equations continuous time
3.2: Physics-informed losses structured objective
3.3: Parameterizing equations ✓/✗ structured model
3.4: Neural operators multi-instance model
3.5: Latent variable model structured models
3.6: System identification ✓/✗ structured model or objective
3.7: Invariance-driven priors struct. model, features, data or objective
4.1: Multi-task learning ✓/✗ ✓/✗ multi-instance data
4.2: Meta learning ✓/✗ ✓/✗ multi-instance data
4.3: Neural processes multi-instance and context data
@@
There are several tasks where such physics-informed models are useful. One is complex forecasting tasks, such as weather prediction, where lots of measurement data are available, and traditional physics models struggle to model such chaotic systems that involve multiple spatial and temporal scales (Schultz et al., 2021; Kurth et al., 2023). Another is solving inverse problems (Ghattas and Willcox, 2021), where a realistic physics model can be examined and queried to calculate what input conditions a system needs for a desired output. These results can then inform the design and operation of the system. Additionally, active fault monitoring is crucial in industry for ensuring the smooth operation of production. By integrating physical models with real-time measurements, these systems assess operational accuracy and predict potential failures, ensuring system reliability (Aldrich and Auret, 2013). This survey covers the diverse ways physics knowledge can be incorporated into ML across datasets ${\mathcal{D}}$, objectives ${\mathcal{L}}$, and models ${\bm{u}}_{\bm{\theta}}$ (Figure 1). Physics-informed models are architectures and prediction techniques informed by domain knowledge, covered in Sections 3.1, 3.3, and 3.5. Learning arbitrary models with physics-informed losses is discussed in Sections 3.2 and 3.6. Various ways to encode physics knowledge in the form of invariances into models are surveyed in section 3.7. Data-driven learning across multiple datasets and experiments includes neural operators (section 3.4), multi-task learning (section 4.1) and meta-learning (section 4.2), and has been used for both physics-informed and -uninformed models and losses. Finally, contextual data is another data-driven approach to incorporate additional domain knowledge, as seen in operator learning (section 3.4) and neural processes (section 4.3). This review is structured to cover many of these cases, as shown in Table 1.
Figure 1: Various ways of incorporating data and machine learning methods. Note that for simplicity, the illustration denotes the solution ${\bm{u}}_{\bm{\theta}}$ as the model of interest, but we may be more broadly interested in learning a family of solutions, the operator, or a simulator of the system with these methods, depending on the data available and what form of predictions are desired from the model.
@@
Machine learning that leverages prior knowledge from sciences and engineering is a rapidly developing field. Several surveys have already been conducted, focusing on specific techniques like physics-informed neural networks (PINNs) (Karniadakis et al., 2021; Pateras et al., 2023), neural operators (Huang et al., 2022a), Koopman operators (Brunton et al., 2022), or exploring how best to combine scientific knowledge with machine learning methods (Von Rueden et al., 2021; Seyyedi et al., 2023). The contribution of this survey is to focus on the challenges of physics-informed learning from the machine learning perspective. This includes looking at physics-informed architectures and loss functions, as well as models and methods such as latent variable models, data augmentation, multi-task learning, and meta-learning that are applicable to the physics-informed setting. The survey is designed to fulfill several key objectives. Firstly, we aim to identify and communicate the outstanding challenges in physics-informed learning to machine learning researchers, providing a pathway for future research directions. Secondly, the survey offers critical guidance, positioning physics-informed learning concepts within the broader landscape of machine learning and connections between various engineering domains. Thirdly, we establish a refined method taxonomy, providing a structured framework for categorizing and distinguishing the different approaches. Additionally, the survey assesses frameworks for evaluating the industrial relevance of these methods and use cases. Lastly, we identify future trends that could shape industrial applications. Through these contributions, the survey seeks to bridge the gap between theoretical research and practical implementation, making physics-informed machine learning more powerful and useful to tackle tough real-world problems.
@@ 2.2 Partial Differential Equations
${\mathsf{L}}_{\bm{a}}$ is a differential operator that contains partial derivatives of the function ${\bm{u}}:\Omega\rightarrow{\mathbb{R}}^{d_{u}}$ w.r.t. spatial coordinates and (potentially also) time, and $\Omega$ is the space-domain domain containing the set of all points in both a bounded space and a time interval over which equation  (3) is defined. ${\bm{f}}({\bm{x}})$ represents a source term that may influence the system’s behavior, and ${\bm{a}}:\Omega\rightarrow{\mathbb{R}}^{d_{a}}$ are the parameters of the differential operator, which could vary across ${\bm{x}}$. Moreover, ${\bm{b}}$ is the boundary condition and ${\bm{n}}$ is normal to the spatial boundary. Notably, equation  (4) addresses the initial conditions at the starting time and accommodates three principal types of boundary conditions, i.e., Dirichlet, Neumann, and Robin boundary conditions depending on the value of ${\bm{c}}_{1}$ and ${\bm{c}}_{2}$.
@@
While the analytical solution of PDEs is restricted to very specific problems, there have been enormous developments in numerical methods for approximating ${\bm{u}}$ and these methods rely on discretization, which is a process that transforms the continuous nature of space and time into a series of discrete points. Spatially, the common approaches are finite difference, finite volume, and finite element methods. Temporally, one breaks the time interval into multiple steps. Here, methods like the method of lines convert the PDE to a system of ordinary differential equations solvable by techniques such as Runge-Kutta. Even a finite difference scheme can be used for temporal discretization. Beyond these separate approaches, advanced methods like finite element space-time methods tackle both spatial and temporal aspects together, offering increased efficiency and accuracy for complex problems. Needless to say, there are many numerical methods developed in approximating ${\bm{u}}$ in a PDE, we only focus on the few common approaches in this review. The finite difference method (FDM) approximates derivatives by differences, essentially replacing continuous changes with discrete steps. It then solves the resulting equation on a grid of points (Thomas, 2013). The finite volume method (FVM), similar to FDM, focuses on the conservation laws over discrete volumes, such as fluid dynamics and heat transfer problems (Eymard et al., 2000). The finite element method (FEM) breaks down the domain into smaller, simpler parts called elements. Within each element, the solution is approximated using specific basis functions that can accurately represent the behavior of $u$ (Szabó and Babuška, 2021). Spectral methods, which employ globally defined high-order basis functions, offer exceptional accuracy for problems with regular geometries. These methods represent the solution as a sum of mathematical functions like Fourier series or polynomials on the domain (Canuto et al., 2007). In contrast, mesh-free methods do not require a pre-defined mesh. They employ scattered points throughout the domain to represent the solution, making them well-suited for problems with complex geometries or discontinuities (Griebel et al., 2005).
@@ 2.3 System Identification
System identification is the practice of building mathematical models of dynamical systems using measurements of their inputs and outputs. This practice is fundamental across science and engineering in order to ground and calibrate mathematical models to the observed reality.
@@
which considers the prediction error for $K$ trajectories over a time horizon $h$. This captures the prediction problem’s sequential nature but requires a more expensive objective that may be harder to optimize. A probabilistic approach considers the joint distribution $p({\bm{Y}},{\bm{X}}\mid{\bm{W}},{\bm{\theta}})$ which, continuing the Markov assumption, factorizes into $p({\bm{x}}_{1})\prod_{t=1}^{T}p({\bm{y}}_{t}\mid{\bm{x}}_{t},{\bm{\theta}})p({\bm{x}}_{t+1}\mid{\bm{x}}_{t},{\bm{w}}_{t},{\bm{\theta}})$. The latent states ${\bm{X}}$ are inferred using Bayesian filtering and smoothing algorithms (Särkkä and Svensson, 2023), and the model parameters ${\bm{\theta}}$ are optimized to maximize the log marginal likelihood of the measurements ${\mathcal{L}}_{\textsc{lml}}({\bm{\theta}})=\log\int p({\bm{Y}},{\bm{X}}\mid{\bm{W}},{\bm{\theta}})\,{\textnormal{d}}{\bm{X}}$ using algorithms such as expectation maximization (EM). Approximate inference techniques are required if the posterior $p({\bm{X}}\mid{\bm{Y}},{\bm{W}},{\bm{\theta}})$ can not be computed in closed form.
@@ 2.4 Function Approximation with Neural Networks
For sequential tasks that require memory, recurrent neural networks have a stateful representation state ${\bm{z}}^{s}$ so that $({\bm{y}}_{t},{\bm{z}}^{s}_{t+1})={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},{\bm{z}}^{s}_{t})$. A non-Markovian alternative is self-attention, which computes its own attention weights that dictate which values inform the internal representation. Mathematically, this is written as ${\bm{Z}}^{(l)}=\text{Softmax}({\bm{Q}}_{\bm{\theta}}({\bm{Z}}^{(l-1)}){\bm{K}}_{\bm{\theta}}({\bm{Z}}^{(l-1)})^{\top}/\sqrt{d}){\bm{V}}_{\bm{\theta}}({\bm{Z}}^{(l-1)})$, where ${\bm{Z}}$ is a sequence of representations. The softmax operation and ‘query’ ${\bm{Q}}$ and ‘key’ ${\bm{K}}$ terms compute the self-attention weights that scale the ‘value’ terms ${\bm{V}}$. The outer product in the softmax means self-attention scales quadratically in the sequence length. Self-attention can be stacked in a scalable fashion to form the transformer architecture, which is capable of learning rich correlations between complex sequences of data, such as natural language (Vaswani et al., 2017).
For structured data that can be defined using a graph with nodes $\mathrm{N}$ and edges $\mathrm{E}$, a graph neural network computes its representations by aggregating across the graph using messaging passing, i.e., ${{\bm{z}}_{i}^{(l)}=\sigma\left(\sum_{j\in\mathrm{N}_{i}}{\bm{f}}^{(l)}_{\bm{\theta}}({\bm{z}}_{i}^{(l-1)},{\bm{z}}_{j}^{(l-1)},\mathrm{E}_{ij})\right)}$, applying the function approximator to neighbouring representations (Bronstein et al., 2021).

### v2

@@ Abstract
This survey examines the broad suite of methods and models for combining machine learning with physics knowledge for prediction and forecasting, with a focus on partial differential equations. These methods have attracted significant interest because of their potential impact on the advancement of scientific research and industrial practices, promising improvements to using small- or large-scale datasets and expressive predictive models with useful inductive biases. The survey has two parts. The first considers incorporating physics knowledge on an architectural level through objective functions, structured predictive models, and data augmentation. The second considers data as physics knowledge, which motivates looking at multi-task, meta, and contextual learning as an alternative approach to incorporating physics knowledge in a data-driven fashion. Finally, we also provide an industrial perspective on the application of these methods and a survey of the open-source ecosystem for physics-informed machine learning.
@@ 1 Introduction
Prediction lies at the heart of science and engineering, and many advances involve the discovery of simple patterns — often expressed as symbolic expressions — that approximate the evolution of our physical world. Still, these mathematical models are only achieved through some degree of simplification and abstraction, and thus rarely capture the complexity of the physical system in its full fidelity. As a result, there is significant interest in learning models from only measurements of the real world, spurring the development of machine learning (ML) (Bishop and Nasrabadi, 2006) for the physical sciences. This field relies on using large datasets of measurements to constrain highly flexible parametric models rather than evoking domain knowledge.
@@
However, for applications that require a high degree of reliability, robustness, and trust in their predictions, purely data-driven models could be deemed insufficient, especially when there is limited data. A natural compromise is to build predictive models that leverage the prior knowledge obtained through centuries of scientific study with the flexibility and scale of modern machine learning (Karniadakis et al., 2021; Karpatne et al., 2022). Achieving this requires models that can integrate prior knowledge in the form of inductive biases (Baxter, 2000), e.g., additional task-informed structure incorporated into the model, objective, or learning algorithm. This added structure acts as a filter, essentially guiding the model to focus on solutions that align with scientific understanding. As a result, the model’s search space becomes smaller, with less irrelevant territory to explore. This targeted exploration allows the model to learn effectively even with limited data. The key challenge lies in designing these inductive biases carefully. If the model is too restrictive, the resulting solution might be suboptimal due to underfitting. This review surveys such methods to inform machine learning models with prior physics knowledge for prediction.
Section Solve the Infer the Generalize Inductive bias
simulation system across systems
3.1: Differential equations prediction in continuous time
3.2: Parameterizing equations physics-informed model
3.3: Physics-informed losses physics-informed objective
3.4: Neural operators multi-instance learning
3.5: Latent variable model latent features and dynamics
3.6: System identification ✓/✗ physics-informed models, objectives
3.7: Model invariances physics-inspired invariances
4.1: Multi-task learning ✓/✗ ✓/✗ multi-instance learning
4.2: Meta learning ✓/✗ ✓/✗ multi-instance learning
4.3: Neural processes multi-instance, contextual learning
@@
There are several tasks where such physics-informed models are useful. One is complex forecasting tasks, such as weather prediction, where lots of measurement data are available, and traditional physics models struggle to model such chaotic systems that involve multiple spatial and temporal scales (Schultz et al., 2021; Kurth et al., 2023). Another is solving inverse problems (Ghattas and Willcox, 2021), where a realistic physics model can be examined and queried to calculate what input conditions a system needs for a desired output. These results can then inform the design and operation of the system. Additionally, active fault monitoring is crucial in industry for ensuring the smooth operation of production. By integrating physical models with real-time measurements, these systems assess operational accuracy and predict potential failures, ensuring system reliability (Aldrich and Auret, 2013). This survey covers the diverse ways physics knowledge can be incorporated into ML across datasets ${\mathcal{D}}$, objectives ${\mathcal{L}}$, and models ${\bm{u}}_{\bm{\theta}}$ (Figure 1). Physics-informed models are architectures and prediction techniques informed by domain knowledge, covered in sections 3.1, 3.2, and 3.5. Learning arbitrary models with physics-informed losses is discussed in sections 3.3 and 3.6. Various ways to encode physics knowledge in the form of model invariances are covered in section 3.7. Data-driven learning across multiple datasets and experiments includes neural operators (section 3.4), multi-task learning (section 4.1) and meta-learning (section 4.2), and has been used for both physics-informed and uninformed models and losses. Finally, contextual data is another data-driven approach to incorporate additional domain knowledge, as seen in operator learning (section 3.4) and neural processes (section 4.3). This review is structured to cover many of these cases, as shown in Table 1.
Figure 1: This survey covers the diverse ways physics knowledge can be incorporated into ML across datasets ${\mathcal{D}}$, objectives ${\mathcal{L}}$ and models ${\bm{u}}_{\bm{\theta}}$. Physics-informed models are architectures and prediction techniques informed by domain knowledge, covered in sections 3.1, 3.2, and 3.5. Learning arbitrary models with physics-informed losses is discussed in sections 3.3 and 3.6. Data augmentation can also incorporate physics knowledge to encode invariances (section 3.7). Data-driven learning across multiple datasets and experiments includes neural operators (section 3.4), multi-task learning (section 4.1) and meta-learning. Finally, contextual data is another data-driven approach to incorporate additional domain knowledge, as seen in operator learning (section 3.4) and neural processes (section 4.3).
@@
Machine learning that leverages prior knowledge from sciences and engineering is a rapidly developing field. Several surveys have already been conducted, focusing on specific techniques such as physics-informed neural networks (PINNs) (Karniadakis et al., 2021; Pateras et al., 2023), neural operators (Huang et al., 2022a), Koopman operators (Brunton et al., 2022), or exploring how best to combine scientific knowledge with machine learning methods (Von Rueden et al., 2021; Seyyedi et al., 2023). Wang and Yu (2021) recently provided a survey that focuses primarily on deep learning models for dynamical systems with a focus on architectures and objectives, and therefore overlaps with our survey in areas such as physics-based equivariances. However, our survey takes a broader machine learning perspective, investigating how prior physics knowledge can be incorporated through not only architectures, but also datasets, and the combination of the two. The survey is designed to fulfill several key objectives. Firstly, our aim is to identify and communicate the outstanding challenges in physics-informed learning to machine learning researchers, providing a pathway for future research directions. Secondly, the survey offers critical guidance, positioning physics-informed learning concepts within the broader landscape of machine learning and connections between various engineering domains. Thirdly, we establish a refined method taxonomy, providing a structured framework for categorizing and distinguishing the different approaches. Additionally, the survey assesses frameworks for evaluating the industrial relevance of these methods and use cases. Lastly, we identify future trends that could shape industrial applications. Through these contributions, the survey seeks to bridge the gap between theoretical research and practical implementation, making physics-informed machine learning more powerful and useful to tackle tough real-world problems.
@@ 2.2 Partial Differential Equations
While ODEs provide a robust framework for modeling time dependent processes, many real-world phenomena involve interdependencies across multiple dimensions. PDEs extend the principles of ODEs by incorporating spatial variability and are used to model systems where the state depends on multiple variables. This section covers the general form of PDEs, including boundary and initial conditions, and discusses various numerical methods for solving them, highlighting their critical role in modeling.
@@
${\mathsf{L}}_{\bm{a}}$ is a differential operator that contains partial derivatives of the function ${\bm{u}}:\Omega\rightarrow{\mathbb{R}}^{d_{u}}$ with respect to spatial coordinates and (potentially also) time, and $\Omega$ is the space-domain domain containing the set of all points in both a bounded space and a time interval over which equation (3) is defined. ${\bm{f}}({\bm{x}})$ represents a source term that may influence the system’s behavior, and ${\bm{a}}:\Omega\rightarrow{\mathbb{R}}^{d_{a}}$ are the parameters of the differential operator, which could vary across ${\bm{x}}$. Moreover, ${\bm{b}}$ is the boundary condition and ${\bm{n}}$ is normal to the spatial boundary. Notably, equation (4) addresses the initial conditions at the starting time and accommodates three principal types of boundary conditions, i.e., Dirichlet, Neumann, and Robin boundary conditions depending on the value of ${\bm{c}}_{1}$ and ${\bm{c}}_{2}$.
@@
While the analytical solution of PDEs is restricted to very specific problems, there have been enormous developments in numerical methods for approximating ${\bm{u}}$ and these methods rely on discretization, which is a process that transforms the continuous nature of space and time into a series of discrete points. Spatially, the common approaches are finite difference (LeVeque, 2007), finite volume (LeVeque, 2002), and finite element methods (Hughes, 2003). Temporally, one breaks the time interval into multiple steps. Here, methods like the method of lines convert the PDE to a system of ordinary differential equations solvable by techniques such as Runge-Kutta. Even a finite difference scheme can be used for temporal discretization. Beyond these separate approaches, advanced methods like finite element space-time methods tackle both spatial and temporal aspects together, offering increased efficiency and accuracy for complex problems. Needless to say, there are many numerical methods developed in approximating ${\bm{u}}$ in a PDE, we only focus on the few common approaches in this review. The finite difference method (FDM) approximates derivatives by differences, essentially replacing continuous changes with discrete steps. It then solves the resulting equation on a grid of points (Thomas, 2013). The finite volume method (FVM) , similar to FDM, focuses on the conservation laws over discrete volumes, such as fluid dynamics and heat transfer problems (Eymard et al., 2000). The finite element method (FEM) breaks down the domain into smaller, simpler parts called elements. Within each element, the solution is approximated using specific basis functions that can accurately represent the behavior of $u$ (Szabó and Babuška, 2021). Spectral methods, which employ globally defined high-order basis functions, offer exceptional accuracy for problems with regular geometries. These methods represent the solution as a sum of mathematical functions like Fourier series or polynomials on the domain (Canuto et al., 2007). In contrast, mesh-free methods do not require a pre-defined mesh. They employ scattered points throughout the domain to represent the solution, making them well-suited for problems with complex geometries or discontinuities (Griebel et al., 2005).
@@ 2.3 System Identification
Solving ODEs and PDEs is sometimes referred to as the forward problem, simulating a physical system to a satisfactory fidelity. The inverse problem, inferring the parameters of a physical system from data, is also of interest (Colton et al., 1990). This setting has many names. System identification is the classical name from the control theory literature (Åström and Eykhoff, 1971), where typically some domain knowledge is assumed and only control input and measurement data is available. In more modern machine learning terminology, model learning or regression can also be used (Bishop and Nasrabadi, 2006), but these terms apply to a wider set of problems that do not necessarily involve timeseries data, domain knowledge or dynamical systems.
@@
this loss is considered as the prediction error ${\mathcal{L}}$ for $K$ trajectories over a time horizon $h$. This captures the prediction problem’s sequential nature but requires a more expensive objective that may be harder to optimize. A probabilistic approach considers the joint distribution $p({\bm{Y}},{\bm{X}}\mid{\bm{W}},{\bm{\theta}})$ which, continuing the Markov assumption, factorizes into $p({\bm{x}}_{1})\prod_{t=1}^{T}p({\bm{y}}_{t}\mid{\bm{x}}_{t},{\bm{\theta}})p({\bm{x}}_{t+1}\mid{\bm{x}}_{t},{\bm{w}}_{t},{\bm{\theta}})$. The latent states ${\bm{X}}$ are inferred using Bayesian filtering and smoothing algorithms (Särkkä and Svensson, 2023), and the model parameters ${\bm{\theta}}$ are optimized to maximize the log marginal likelihood of the measurements ${\mathcal{L}}_{\textsc{lml}}({\bm{\theta}})=\log\int p({\bm{Y}},{\bm{X}}\mid{\bm{W}},{\bm{\theta}})\,{\textnormal{d}}{\bm{X}}$ using algorithms such as expectation maximization (EM). Approximate inference techniques are required if the posterior $p({\bm{X}}\mid{\bm{Y}},{\bm{W}},{\bm{\theta}})$ can not be computed in closed form.
@@ 2.4 Function Approximation with Neural Networks
For sequential tasks that require memory, recurrent neural networks (RNNs) (Rumelhart et al., 1985) have a stateful representation state ${\bm{z}}^{s}$ so that $({\bm{y}}_{t},{\bm{z}}^{s}_{t+1})={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},{\bm{z}}^{s}_{t})$. A non-Markovian alternative is self-attention, which computes its own attention weights that dictate which values inform the internal representation. Mathematically, this is written as ${\bm{Z}}^{(l)}=\text{Softmax}({\bm{Q}}_{\bm{\theta}}({\bm{Z}}^{(l-1)}){\bm{K}}_{\bm{\theta}}({\bm{Z}}^{(l-1)})^{\top}/\sqrt{d}){\bm{V}}_{\bm{\theta}}({\bm{Z}}^{(l-1)})$, where ${\bm{Z}}$ is a sequence of representations. The softmax operation and ‘query’ ${\bm{Q}}$ and ‘key’ ${\bm{K}}$ terms compute the self-attention weights that scale the ‘value’ terms ${\bm{V}}$. The outer product in the softmax means self-attention scales quadratically in the sequence length. Self-attention can be stacked in a scalable fashion to form the transformer architecture, which is capable of learning rich correlations between complex sequences of data, such as natural language (Vaswani et al., 2017).
For structured data that can be defined using a graph with nodes $\mathrm{N}$ and edges $\mathrm{E}$, a graph neural network computes its representations by aggregating across the graph using messaging passing, i.e., ${{\bm{z}}_{i}^{(l)}=\sigma\left(\sum_{j\in\mathrm{N}_{i}}{\bm{f}}^{(l)}_{\bm{\theta}}({\bm{z}}_{i}^{(l-1)},{\bm{z}}_{j}^{(l-1)},\mathrm{E}_{ij})\right)}$, applying the function approximator to neighboring representations (Bronstein et al., 2021).